# Do Pre-Trained Language Models Detect and Understand Semantic Underspecification? Ask the DUST!

### v1

@@ 1 Introduction
Figure 1: In experiment 1 (top), we test if LMs distinguish underspecified sentences (S1) from more specified counterparts (S2) when explicitly prompted. In experiment 2 (bottom), we test if LMs interpret correctly S1 and S2 in a more naturalistic communicative setting.
@@
However, handling semantically underspecified language is critical for LMs. Since underspecified sentences can license multiple interpretations, choosing one arbitrarily can lead to undesired or even harmful consequences for communication (see Hutchinson et al., 2022; Pezzelle, 2023, for an in-depth discussion of this issue). Thus, LMs should (1) recognize semantically underspecified inputs and (2) interpret them appropriately, ideally with no biases toward a default reading.
@@ 2.1 Semantic Underspecification
Egg (2010) gives a detailed categorization of underspecification, grouping instances into four types based on whether the instance’s constituent parts comprise the same semantic value across its readings, and whether it is possible to give a single syntactic analysis for all the readings.
@@ 3 Dataset
The third type consists of syntactically but not semantically homogeneous expressions, which share a single syntactic structure but do not share the same semantic material in its constituent parts. It is unique in that, besides instances of referential ambiguity, all examples of type 3 are underspecified but not ambiguous. It is thus particularly interesting for our work, as underspecified but not ambiguous expressions are important for LMs to handle, but currently understudied. Examples include deictic expressions and expressions that are underspecified due to missing information.
@@
Egg’s fourth and final type, consisting of expressions that are neither syntactically nor semantically homogeneous contains homonymic expressions. These contain a word with multiple meanings; the chosen reading of the homonym can change the meaning of the whole expression. We exclude this type of underspecification from DUST, as it applies mainly to words and phrases. While a word (e.g. “plant”) can have multiple syntactically and semantically distinct readings, most sentences that contain homonyms have readings that rely on the same syntactic structure. For example, “he walked to the bank” contains the homonym “bank”, which could refer to a riverbank of a financial institution, but the syntactic structure of the sentence is the same across both readings. In Appendix A, we report the results and limitations of a preliminary attempt to construct examples falling into this type.
@@ 5 Detecting Semantic Underspecification
We test whether LMs recognize that a sentence is more or less specified by explicitly asking them via a prompt. This task seeks to replicate the process of assessing human speakers’ understanding of this phenomenon by asking them to provide a metalinguistic assessment.
@@ 5.1 Experimental Setup
should be judged by LMs as more likely than the input where the blue (underspecified) and red (more specified) sentences are switched.
To ensure prompt diversity, and because models may not have been exposed to terminology such as “underspecified” during training, we also use alternate versions of our prompts, where “under-” and “more specified” are replaced by “(un)ambiguous’, or “contain (little/a lot of) (information/detail)”. We also create prompts that reverse the order in which the under- and more specified sentences are presented. [^4] For more example inputs, see Appendix D. In total, we create 33,968 input pairs: 2,123 sentence pairs $\times$ 4 prompt variants $\times$ 4 orders. Then, for each model and input pair, we record whether the model correctly assigns a lower perplexity to the specification-matched input than to its mismatched counterpart.
Figure 2: Proportion of specification-matched inputs that receive lower perplexity than their specification-mismatched counterparts; higher is better. Asterisks (*) indicate performance significantly ($p<0.05$) different from chance (50%). The leftmost column, Overall, reports the proportion computed over the whole dataset.
@@
Our experimental setup aims to determine if models can recognize underspecification when prompted to do so; however, prior work suggests that models may struggle to understand their prompts Webson and Pavlick (2022). So, we first verify the soundness of our setup by using it to test models in an easier domain: sentiment. We gather “very positive” and “very negative” sentences from the SST-5 dataset Socher et al. (2013), and insert them into prompts of the form “This is a positive sentence: [sentence1]. This is a negative sentence: [sentence2]”. All models assign lower perplexity to sentiment-matched inputs with at least 65% and an average of 75% accuracy (see Appendix F for details). This indicates that current LMs can be tested using this experimental setup.
Type Sentence OPT-13B Llama2-7B Mistral 7B
1 Danny left Yevgeni; also Andrei 0.625 0.8125 0.875
Yevgeni approached Andrei; also Danny 0.6875 0.6875 0.8125
2 Andrei looked at Danny moving a blue telescope 0.5 0.625 0.8126
Andrei left the person with a green chair 0.5 0.625 0.6875
3 Pay the fee 0.6875 0.5 0.875
Do this with force, and the can should be smashed 0.4375 0.5 0.5
Table 2: Underspecified sentences and proportion of examples (across prompts and orderings) where the specification-matched input received lower perplexity than its mismatched counterpart; higher is better.
@@ 5.2 Results
Our results (Figure 2) indicate that some LMs can recognize underspecification. All models besides Flan-T5 do so at a rate significantly higher than chance; Flan-T5’s poor performance may be due to its architecture, i.e., a fine-tuned encoder-decoder model and not a decoder-only model trained on causal language modeling like the others. Stronger models more often prefer specification-matched prompts: Mistral, the best-performing model across the board, achieves an overall accuracy of 0.74. The second-best model, Llama 2, lags behind Mistral by almost 10 accuracy points, with an overall accuracy of 0.65. In turn, this LM outperforms OPT (0.55) and GPT-2 (0.53) by as many accuracy points, indicating a clear ranking between the various models.
@@
To shed light on the cases where each model succeeds and fails, we conduct a qualitative analysis on a handful of samples, i.e., two underspecified sentences per DUST type, with the best-performing Mistral, Llama 2, and OPT LMs. As reported in Table 2, Mistral is consistently the best-performing model in these examples—it achieves the highest proportion of correctly assigned perplexities. Moreover, the numbers closely align with the patterns reported in Figure 2: for both Mistral and Llama 2, type 1 is the easiest, followed by types 2 and 3. Taking a closer look at the examples, it is worth noting that, (1) for “Pay the fee”, Llama 2 does not recognize it as more underspecified than its counterpart “Pay the licensing fee” (not reported in the table), while OPT and even more Mistral do; (2) for “Do this with force, and the can should be smashed”, all models similarly struggle; none of them recognize the sentence “Do this part with force, and the can should be smashed” (not reported) as its more specified counterpart.
@@
Overall, our results suggest that modern LMs can moderately identify underspecification if explicitly asked to do so via a prompt. However, our current approach does involve significant prompting, which has several disadvantages when evaluating LMs’ linguistic abilities. Model performance is highly sensitive to the specific prompt used Mizrahi et al. (2024). Moreover, previous work has shown that using meta-linguistic prompts, which explicitly ask for a model’s linguistic judgment, often yields different results than evaluating linguistic tasks using naturalistic data Hu and Levy (2023). While our experiments do not directly ask models if a given sentence is underspecified (we instead compare two versions of the same sentence), our inputs still do not reflect naturalistic data. Our second experiment is motivated by the need to account for this issue.
@@ 6.1 Experimental Setup
[…]
@@
We then embed these sentences and continuations in simple templates like “[sentence]. That is, [continuation]”. Given an underspecified sentence, where both continuations are equally plausible, LMs should ideally assign roughly equal perplexity to both. [^6] Underspecified sentences could still have a more frequent “default reading”, making one continuation more likely. For more specified sentences, in contrast, only one interpretation is possible; the corresponding continuation should therefore receive more probability than the incompatible one.
@@ 6.2 Results
Figure 4: Proportion of inputs where the plausible continuation to a specified sentence received lower perplexity; higher is better. Asterisks (*) indicate performance that differs significantly ($p<0.05$) from chance (50%).
@@
Our results (Figure 4) indicate that the type of prompt used heavily influenced the degree to which models preferred the correct, plausible continuation; however, unlike in experiment 1, performance is poor overall. In the base case, where no prompts were used, only Llama 2 7B performs significantly above chance. With the “that is” prompt, model performance improves, though there is no clear trend in which models perform best. It is only with the most extensive prompt that we recover both better performance and the trend from experiment 1, where newer models perform better.
@@ 7 Discussion
While the present work performs an in-depth evaluation of how LMs behave when faced with semantic underspecification, our research does not explore the inner mechanisms that underlie this capability. We acknowledge that doing so would provide complementary evidence that may be needed to shed full light on the phenomenon.
@@ Appendix A Type 4 of underspecification
In an attempt to collect sentences containing this type of underspecification, we collect sentences containing homonymic expressions from a sample of English Wikipedia (Mukherjee and Bhattacharyya, 2012). We do so by selecting sentences that contain any homonym from a list of 100 homonyms, selected by Maciejewski and Klepousniotou (2016) based on linguistic principles, dictionary entries, and subjective ratings. We created more specified counterparts by selecting random sentences from the that contain none of the homonyms from the list. We collect a total of 980 sentence pairs; note that unlike other pairs in DUST, these are not minimal pairs. One (partial) sentence of this type is
@@
Figure 5: Proportion of specification-matched inputs with lower perplexity than their specification-mismatched counterparts; higher is better. Asterisks (*) indicate performance significantly ($p<0.05$) above or below chance (50%).
@@ Appendix C Model and Experimental Details
Appendix D Experiment 1 Prompts
@@
Appendix E Experiment 2 Prompts
Suppose we have the underspecified sentence ‘Andrei looked at Danny moving a green bag’ and the more specified counterpart ‘Andrei looked at Danny who was moving a green bag ’. Examples of specification-matched prompts we would then obtain, from low to high level of prompting, are:
@@ Appendix E Experiment 2 Prompts
Appendix F Sentiment perplexity
To test whether the experimental design of experiment 1 functions correctly, we first ran this experiment with sentiment classification instead of recognition of underspecification as a goal. The models were prompted with prompts of the form “prompt1: ’sentence1’. prompt2: ’sentence2’.", where the prompts are of the form "This is a positive/negative sentence" and the sentences are sentences rated ‘very positive’ or ‘very negative’ in the SST-5 dataset (Socher et al., 2013). The results of this can be seen in Figure 6.
@@ Appendix F Sentiment perplexity
Appendix G Could models be detecting surface statistics instead of underspecification?

### v2

@@ 1 Introduction
Figure 1: In experiment 1 (top), we test if LMs distinguish underspecified sentences (S1) from more specified counterparts (S2) when explicitly prompted. In experiment 2 (bottom), we test if LMs correctly interpret S1 and S2 in a more naturalistic communicative setting.
@@
However, handling semantically underspecified language is critical for LMs. Since underspecified sentences can license multiple interpretations, choosing one arbitrarily can lead to undesired or even harmful consequences for communication (see Hutchinson et al., 2022; Pezzelle, 2023, for an in-depth discussion of this issue). Correctly processing underspecified language could have real-world consequences for NLP systems. For example, if an embodied virtual assistant were to misinterpret a question like “Can I hang the painting with the cup?”, there could be risks: while the answer may be yes if the cup is the subject of the painting, the system should answer no if this is not the case. In machine translation, models should carefully handle underspecification; for example, when translating “they”, models must determine whether the pronoun refers to a group of people or an individual of unknown, nonbinary, or intentionally underspecified gender. Therefore, LMs should (1) recognize semantically underspecified inputs and (2) interpret them appropriately, ideally with no biases toward a default reading.
@@ 2.1 Semantic Underspecification
Egg (2010) gives a detailed categorization of underspecification, grouping instances into four types based on whether the instance’s constituent parts comprise the same semantic value across its readings, and whether it is possible to give a single syntactic analysis for all the readings. As this classification allows us to explore the interaction between the semantic and syntactic dimensions of underspecification and their effects on LMs, we use it as the theoretical backbone to build our dataset.
@@ 3 Dataset
The third type consists of syntactically but not semantically homogeneous expressions, which share a single syntactic structure but do not share the same semantic material in its constituent parts. It is unique in that, besides instances of referential ambiguity, all examples of type 3 in DUST are underspecified but not ambiguous. It is thus particularly interesting for our work, as underspecified but not ambiguous expressions are important for LMs to handle, but currently understudied. Examples include deictic expressions and expressions that are underspecified due to missing information.
@@
Egg’s fourth and final type, consisting of expressions that are neither syntactically nor semantically homogeneous, concerns phenomena that occur at the word level, such as homonymic expressions. While a word (e.g. “plant”) can have multiple syntactically and semantically distinct readings, most sentences that contain homonyms have readings that rely on the same syntactic structure. For example, “he walked to the bank” contains the homonym “bank”, which could refer to a riverbank of a financial institution, but the syntactic structure of the sentence is the same across both readings.
Unlike the three types described above, the phenomena belonging to this type cannot be easily studied using an experimental setup based on minimal pairs—the one used in this work—where the two interpretations are embedded in two phrases or sentences that only differ by a single, minimal intervention (see Appendix A for a preliminary exploration of the problem). For this reason, we do not include type 4 in our benchmark and leave a comprehensive exploration of it to future work.
@@ 4 Models
Except for Flan-T5, the models we use in our research were not instruction-tuned during pre-training. This allows us to test the abilities of base LMs to process underspecification.
@@ 5 Detecting Semantic Underspecification
We test whether LMs recognize that a sentence is more or less specified by looking at the perplexity it assigns to a prompt comparing the degree of (under)specification of two embedded sentences. This task seeks to replicate the process of assessing human speakers’ understanding of this phenomenon by asking them to provide a metalinguistic assessment. We use a perplexity-based approach, rather than evaluating the generative behavior of the models in response to a prompt, as there is growing evidence that prompt-based approaches are not suitable for this purpose Hu and Levy (2023). For comparison, we include a preliminary experiment exploring model-generated responses in Appendix D.
@@ 5.1 Experimental Setup
should be judged by LMs as more likely than the input where the blue (underspecified) and red (more specified) sentences are switched. We test this by computing, for a given prompt, the product of the model perplexities assigned to each token in it.
To ensure prompt diversity, and because models may not have been exposed to terminology such as “underspecified” during training, we also use alternate versions of our prompts, where “under-” and “more specified” are replaced by “(un)ambiguous’, or “contain (little/a lot of) (information/detail)”. We also create prompts that reverse the order in which the under- and more specified sentences are presented. [^4] For more example inputs, see Appendix E. In total, we create 33,968 input pairs: 2,123 sentence pairs $\times$ 4 prompt variants $\times$ 4 orders. Then, for each model and input pair, we record whether the model correctly assigns a lower perplexity to the specification-matched input than to its mismatched counterpart.
[…]
@@
Our experimental setup aims to determine if models can recognize underspecification when prompted to do so; however, prior work suggests that models may struggle to understand their prompts Webson and Pavlick (2022). So, we first verify the soundness of our setup by using it to test models in an easier domain: sentiment. We gather “very positive” and “very negative” sentences from the SST-5 dataset Socher et al. (2013), and insert them into prompts of the form “This is a positive sentence: [sentence1]. This is a negative sentence: [sentence2]”. All models assign lower perplexity to sentiment-matched inputs with at least 65% and an average of 75% accuracy (see Appendix G for details). This indicates that current LMs can be tested using this experimental setup.
Phenomenon Sentence OPT Llama Mistral
Logical Form Danny approached Andrei; Yevgeni, too 0.25 0.75 1
Ellipsis Andrei and Danny put down a yellow chair 0.5 0.75 1
PP attach. amb. Andrei approached the person with a yellow bag 0.5 0.75 1
VP attach. amb. Danny looked at Andrei moving a green chair 0.5 0.75 1
Conjunction amb. Andrei and Danny held the yellow bag and chair 0.75 0.75 1
Referential amb. Although they ran at about the same speed, 0.25 0.25 0.25
Sue beat Sally because she had such a bad start
Added compound Get yourself a flannel shirt and wear it 0.75 0.75 0.75
over a plain tee shirt.
Fused head This means you have broken the seal and 0.5 1 0.75
can now twist off the lid.
Implicit reference 4. Do not slurp. 0.75 0.75 1
Metonymic ref. Think about your plant’s activity 0.25 1 1
Table 2: Underspecified sentences and proportion of examples (across prompts and orderings) where the specification-matched input received lower perplexity than its mismatched counterpart for OPT-13B, Llama2-7B, and Mistral 7B; higher is better. Horizontal lines demarcate the types of underspecification as per Egg’s (2010) taxonomy. Phenomena containing underspecification without ambiguity are italicized. ‘amb.’ stands for ambiguity, ‘ref.’ for reference.
@@ 5.2 Results
[…]
@@
To shed light on the cases where each model succeeds and fails, we conduct a qualitative analysis on a handful of samples, i.e., one underspecified sentence per linguistic phenomenon, with the best-performing Mistral, Llama 2, and OPT LMs. The actual sentences considered in this analysis can be found in Table 2.
Among types 1 and 2 of underspecification, we find that for almost all phenomena, the qualitative inspection closely mirrors the quantitative results reported in  Figure 2 – Mistral is consistently better than all other models, and Llama 2 outperforms OPT. However, this is much less the case for the conjunction scopal ambiguity, where both OPT-13B and Llama2-7B perform at a similar level and are much closer to Mistral. This is in line with the overall quantitative trends in Appendix F. We hypothesize that this may be because the underspecified sentences displaying this phenomenon can be considered difficult to parse. If true, this would suggest that LMs may use some notion of sentence complexity (e.g. the difficulty to parse it) as a stand-in for underspecification.
We also observe that, for referential ambiguity, performance for all models is very low. This may be because most sentences containing referential ambiguity are part of the Winograd Schema Challenge dataset, which may be part of the training data of the tested LMs. As an effect of this, the perplexity assigned to these underspecified sentences may be lower than that of the more uncommon control sentences.
@@
Overall, our results suggest that modern LMs can moderately identify underspecification if explicitly asked to do so via prompting. In particular, the observation that newer models perform better suggests that pre-training with more and better data, using more parameters, and relying on various (even if minor) architectural improvements could eventually lead to models that can accurately recognize underspecified language.
However, our current approach does involve significant prompting, which has several disadvantages when evaluating LMs’ linguistic abilities. Model performance is highly sensitive to the specific prompt used Mizrahi et al. (2024). Moreover, previous work has shown that using meta-linguistic prompts, which explicitly ask for a model’s linguistic judgment, often yields different results than evaluating linguistic tasks using naturalistic data Hu and Levy (2023). While our experiments do not directly ask models if a given sentence is underspecified (we instead compare two versions of the same sentence), our inputs still do not reflect naturalistic data. Our second experiment is motivated by the need to account for this issue.
@@ 6.1 Experimental Setup
We create two specified versions of each DUST sentence originating from the LAVA dataset, corresponding to the possible readings of the original sentence. We also create two continuations to the sentence, which again correspond to distinct readings of the original. Each continuation is compatible with the underspecified sentence, but only one of the fully specified sentences. We slightly adjust the LAVA sentences to make them more correct. [^5] Uses of ‘put-down’ and ‘picked-up’ were changed to ‘put down’ and ‘picked up’, and sentences of the form “NNP1 V NNP2. Also NNP3.” were changed to “NNP1 V NNP2; NNP3 too.”
@@
We then embed these sentences and continuations in simple templates like “[sentence]. That is, [continuation]”. Given an underspecified sentence, where both continuations are equally plausible, LMs should ideally assign roughly equal perplexity to both. [^6] Underspecified sentences could still have a more frequent “default reading”, making one continuation more likely in human speakers’ interpretation. For example, when resolving referential ambiguity, it has been documented that a reader might default to the first possible referent encountered in the text. Alternatively, it has been proposed that, when parsing a sentence, speakers could leave ambiguities unresolved if resolving is not strictly necessary Swets et al. (2008). Since we consider both sentence readings in our experiment, we hypothesize that the effects of such default readings cancel out and therefore do not affect our results. For more specified sentences, in contrast, only one interpretation is possible; the corresponding continuation should therefore receive a higher probability than the incompatible one.
@@ 6.2 Results
[…]
@@
[…]
@@
This noisiness is also reflected in the differences between linguistic phenomena and between type of prompt. We find that for referential and VP attachment ambiguity, a greater proportion of inputs resulted in the correct continuation receiving lower perplexity. This seems to suggest that correctly interpreting sentences containing these two phenomena is handled better by the models, although this differs greatly per model and type of prompt. However, we note that relatively few sentences contain these linguistic phenomena, limiting our ability to draw strong conclusions from this observation.
@@ 7 Discussion
While the present work performs an in-depth evaluation of how LMs behave when faced with semantic underspecification, our research does not explore the inner mechanisms that underlie this capability. We acknowledge that doing so would provide complementary evidence that may be needed to shed full light on the phenomenon. Moreover, research could focus on how LMs handle underspecification in more naturalistic scenarios, e.g., in the context of real-world NLP applications, which is something the current work does not explore.
@@
Acknowledgements
Some of the contents of this work are based on FW’s Master’s thesis. Full acknowledgments from the first author can be found therein. We thank the anonymous ARR reviewers for their valuable feedback and the members of the Dialogue Modelling Group at the University of Amsterdam for their insightful comments. MH’s research is supported by an OpenAI Superalignment Fellowship.
@@ Appendix A Type 4 of underspecification
In an attempt to collect sentences containing this type of underspecification, we collect sentences containing homonymic expressions from a sample of English Wikipedia (Mukherjee and Bhattacharyya, 2012). We do so by selecting sentences that contain any homonym from a list of 100 homonyms, selected by Maciejewski and Klepousniotou (2016) based on linguistic principles, dictionary entries, and subjective ratings. We created more specified counterparts by selecting random sentences from the that contain none of the homonyms from the list. We collect a total of 980 sentence pairs; note that unlike other pairs in DUST, these are not minimal pairs. One (partial) sentence of this type (though not from our dataset) is
@@
[…]
@@ Appendix C Model and Experimental Details
Appendix D MCQ experiment
In this experiment, we test whether language models can recognize semantic underspecification by explicitly asking them to generate an answer about the underspecification of a sentence pair. In practice, we prompt GPT2-xl, OPT-13B, and Llama2-7B to generate a response using the following prompt:
Here are two sentences. A: ‘Andrei left the chair with a green telescope’. B: ’Andrei left the chair on which lay a telescope’. Which one of these is more semantically underspecified? Please respond by outputting only A or B. Answer:
where the red and blue sentences are replaced by either an underspecified or its corresponding control sentence from the dataset. The order in which the two sentences are placed is randomized to prevent bias in the model from unduly influencing the final accuracy. Each sentence pair in the dataset is included in a prompt once. In  Table 4, we report model accuracy and number of A, B, or other responses.
model acc. #A #B #other
GPT2-xl 0.31 696 638 789
OPT-13B 0.49 2105 18 0
Llama2-7B 0.48 888 1138 97
Table 4: MCQ task. Model accuracy and number of responses per class by GPT2, OPT, and Llama 2.
The results show that the models perform very poorly when the task is formulated as a multiple-choice task. While the low accuracy might be a result of the models being unable to do this task—something that could perhaps improve when using instruction-tuned models—the observation that all the models are either biased towards one of the options or unable to consistently answer the prompt with one of the two given options, or both, suggests that they are incapable of performing the task in this experimental setup. This matches earlier findings (e.g., Hu and Levy, 2023) and validates our decision to perform perplexity-based evaluation over a MCQ-like type of experiment even further.
Appendix E Experiment 1 Prompts
@@
Appendix F Experiment 1 Results per Phenomenon
In Figure 6, we report the results of Experiment 1 split by linguistic phenomenon.
[…]
Appendix G Sentiment perplexity
To test whether the experimental design of experiment 1 functions correctly, we first ran this experiment with sentiment classification instead of recognition of underspecification as a goal. The models were prompted with prompts of the form “prompt1: ‘sentence1’. prompt2: ‘sentence2’.", where the prompts are of the form "This is a positive/negative sentence" and the sentences are sentences rated ‘very positive’ or ‘very negative’ in the SST-5 dataset (Socher et al., 2013). The results of this can be seen in Figure 7.
Appendix H Experiment 2 Prompts
Suppose we have the underspecified sentence ‘Andrei looked at Danny moving a green bag’ and the more specified counterpart ‘Andrei looked at Danny who was moving a green bag ’. Examples of specification-matched prompts we would then obtain, from low to high levels of prompting, are:
@@ Appendix H Experiment 2 Prompts
Andrei looked at Danny who was moving a green bag. That is, Danny was moving a green bag.
@@
Andrei looked at Danny who was moving a green bag. That is, Andrei was moving a green bag.
@@
Appendix I Could models be detecting surface statistics instead of underspecification?